OpenStudy (anonymous):

Tanner has b more bouncy balls than Hiro. Hiro has 52 bouncy balls. Choose the expression that shows how many bouncy balls Tanner has.

OpenStudy (ksaimouli):

i think b -52=bouncyballs

OpenStudy (anonymous):

T = H + b T = 52 + b

OpenStudy (anonymous):

f(x)=5x^2-7x=4, find -13f(7)-6f(-10)